I have trouble with these remaining problems i could use some help answering these thank you
Geometric sequencing
three of the five variables a1, an, r, n, and Sn are given. Find the two variables not given.
a1 = 10, r = 3, n = 5
a1 = 4, an = 324, r = 3
an = 24, r = , n = 4
Sn = 1022, r = 2, n = 9

Answer by venugopalramana(3286)

I have trouble with these remaining problems i could use some help answering these thank you
Geometric sequencing
three of the five variables a1, an, r, n, and Sn are given. Find the two variables not given.
FORMULAE ARE
AN=A1*R^(N-1)
SN=A1*(R^N -1)/(R-1)
a1 = 10, r = 3, n = 5......AN=10*3^(5-1)=810...SN =10*(3^5-1)/(3-1)=1210.
a1 = 4, an = 324, r = 3....324=4*3^(N-1)...N=5...SN=4*(3^5 -1)/(3-1)=484
an = 24, r = 2, n = 4....24=A1*2^(4-1)...A1=3...SN=3*(2^4 -1)/(2-1)=45
Sn = 1022, r = 2, n = 9 ...1022=A1(2^9 -1)/(2-1)=A1*(512-1)=511A1...A1=2
AN=2*2(9-1)=512